Here's another one...*smiles*
"Egg nog is made by mixing eggs, cream, milk and sugar together. The recipe calls for 3% butterfat milk. Whole milk contains 4% butterfat and skim milk contains 0% butterfat. How much skim milk should be added to 32oz of whole milk to obtain a mixture that is 3% butterfat?"
I got 10.666666666666667 ounces...?
And then the table asks us to list the amount in ounces (whole milk: 32oz...skim milk: x...mixture: x+32) and the amount of non fat for the 96% non fat (whole milk), the 100% non fat (skim milk) and the 97% non fat (mixture)...
What do they mean list the amount of non fat...? Do they mean in equation form like...
whole milk is .04*32=1.28
So 32oz-1.28= 30.72oz non fat...?
Thanks in advance for all of your help! *smiles*

Answer by ptaylor(2048)

Let x=amount of skim milk that needs to be added to the whole milk
Now we know that the amount of pure butterfat that exists before the mixture takes place (0.04*32) has to equal the amount of pure butterfat that exists after the mixture takes place (0.03(32+x)). So our equation to solve is:
0.04*32=0.03(32+x) simplify
1.28=0.96+0.03x
0.03x=1.28-0.96=0.32
x=10.666666666666666666666666667
SAME AS YOUR SOLUTION
GOOOOOOOOD WORK!!!!!!!!!!!!!!!!!

OOOOOOOOOOOOOOOOPs there's more:
Whole Milk---32 oz
Skim milk-----10.66666666667 oz
Mixture-------42.6666666667 oz
What do they mean list the amount of non fat...? Do they mean in equation form like...
I believe that they want you to list the amount (in terms of ounces)of the non-fat components just like you are doing.
Hope this helps----ptaylor